OpenStudy (anonymous):

From an external point E, Tangents ES and ET are drawn to circle O. If measurement of angle E= 78 find the measure of Minor arc ST

OpenStudy (anonymous):



OpenStudy (anonymous):

E = 78, then O = 180 - 78 = 102 deg Perimeter of circle: 2Pi*R. Length of minority arc x: x/2Pi*R = 102/360 --> x = 204*Pi*R/360

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

Angle = Total 360 deg minus <E(78), minus two right angles S and T. < SOT = 360 - (78) - (90) - (90) = 102 deg The ratio: min. length ST/circle perimeter is equal to the ratio: 102 deg/360 deg. Equation: x/2Pi*R = 102/360 x = 102 (2Pi*R)/360 = 204Pi*R/360. When Radius R is given, then we can compute length x.